# Scalaron dark matter dynamics: effects of Higgs non-minimal coupling to gravity

### v1

@@ 4.2 Constraints on scalaron decay to photons
The viability of the scalaron as a DM candidate hinges on its lifetime to be comparable to the age of the universe. Since the scalaron universally couples to the SM matter fields, it can decay to most of the SM particles, if allowed by kinematics. To note, for $m<2m_{e}$, $m_{e}$ being the electron mass, the only possible decay is to a pair of photons, whose rate should be such that the scalaron becomes cosmologically stable. Although the scalaron does not couple to massless gauge fields at the tree-level due to Weyl invariance, [^3] The photon action, $\int d^{4}x\sqrt{-g}~F_{\mu\nu}F^{\mu\nu}$ is invariant under trnasformation defined in Eq.4 the coupling is generated at one loop-level, primarily due to interactions with fermions and massive gauge bosons, as follows Goldberger:2007zk ; Cembranos:2008gj ; Katsuragawa:2017wge :
@@
where $c_{EM}=11/3$ Cembranos:2008gj , for scalaron being lighter than the SM fermions and gauge bosons. Evidently, the bound on the lifetime of DM, i.e. $\tau_{DM}\gtrsim 5\times 10^{18}s$ Audren:2014bca is easily satisfied by the scalaron. However, the photon flux produced from scalaron decays can constrain the decay rate from the observation of an excess photon flux over known astrophysical backgrounds. In this case, gamma-ray observations with SPI, a spectrometer telescope aboard the INTEGRAL satellite, operating in the energy range of 18 keV to 8 MeV Winkler:2003nn will be most sensitive. The photon signal from non-relativistic decays of scalaron in the present epoch yields a monochromatic spectrum, peaking at a single energy bin around, $m/2$. Consequently, the scalaron mass is constrained from the INTEGRAL/SPI data, unlike scenarios with axion-like particles (ALPs) with an undetermined U(1) symmetry breaking scale involved Alonso-Alvarez:2019ssa ; Arias:2012az . Nevertheless, there are existing limits on the decay rate of light DM decaying to two photons Calore:2022pks ; Fischer:2022pse , which can be used to constrain our scalaron scenario. In the right panel of Fig.4, we show the exclusion region on the decay rate (green shaded region) using the INTEGRAL/SPI data, taken from Ref.Calore:2022pks . Authors of Ref.Calore:2022pks analyzed 16 years of the INTEGRAL/SPI data, taking the NFW profile of galactic DM density to constrain the decay rate at 95% confidence level. We compare the decay rate of the scalaron (shown in red solid line) as derived in Eq.32 to find the upper bound of the scalaron mass as $m_{\rm max}=0.7$ MeV. This bound is derived with the assumption that the scalaron satisfies total DM density, otherwise the exclusion limits become weaker.
5 Summary and Outlook
@@
To summarize, we have studied the scalaron DM dynamics in presence of scalaron-Higgs trilinear interaction, which consists of Higgs quartic coupling and Higgs non-minimal coupling to gravity. Depending on the relative contributions of these couplings the scalaron dynamics is modified and its mass is significantly altered to account for observed DM density. In particular, we provide a natural explanation for vanishing scalron-Higgs interaction, leading to keV-MeV DM candidate.
@@
We note that the interplay between Higgs quartic coupling and the non-minimal coupling in the light DM context is demonstrated with a possibility of detection in highly sensitive X-ray/gamma ray telescopes. However, there is significant uncertainty over the experimental determination of these couplings, which can affect scalaron dynamics in the early universe. In fact, measuring the Higgs quartic coupling at colliders is challenging, as it requires cross-section measurements of processes involving the three-Higgs vertex, which are highly suppressed Fuks:2017zkg ; Stylianou:2023tgg . On the other hand, the limits on the Higgs non-minimal coupling in low energy effective theory with $R^{2}$ modification to the Einstein’s GR is an unexplored topic, to the best our knowledge. Here, we have used the LHC data to give an upper bound on the non-minimal coupling assuming the $R^{2}$ term subsides for in the late universe. Therefore, the scalaron dynamics and its mass limits can modify with improvements in understanding of these Higgs interaction along with its experimental verifications.
@@ Acknowledgment
SGC acknowledges funding from the Anusandhan National Research Foundation (ANRF), Govt. of India, under the National Post-Doctoral Fellowship (File no.PDF/2023/002066). SGC thanks Ananda Dasgupta for useful discussions. DG acknowledges the Post-doctoral Fellowship from IISER Kolkata for most of the duration of this work, and from SINP Kolkata for the remaining period.

### v2

@@ 4.2 Constraints on the scalaron decay to photons
The viability of the scalaron as a DM candidate hinges on its lifetime to be comparable to the age of the universe. Since the scalaron universally couples to the SM matter fields, it can decay to most of the SM particles, if allowed by kinematics. To note, for $m<2m_{e}$, $m_{e}$ being the electron mass, the only possible decay is to a pair of photons, whose rate should be such that the scalaron becomes cosmologically stable. Although the scalaron does not couple to massless gauge fields at the tree-level due to Weyl invariance, [^3] The photon action, $\int d^{4}x\sqrt{-g}~F_{\mu\nu}F^{\mu\nu}$ is invariant under transformation defined in Eq.4 the coupling is generated at one loop-level, primarily due to interactions with fermions and massive gauge bosons, as follows Goldberger:2007zk ; Cembranos:2008gj ; Shtanov:2025nue ; Katsuragawa:2017wge :
@@
where $c_{EM}=11/3$ Cembranos:2008gj , for scalaron being lighter than the SM fermions and gauge bosons. Evidently, the bound on the lifetime of DM, i.e., $\tau_{DM}\gtrsim 5\times 10^{18}s$ Audren:2014bca is easily satisfied by the scalaron. However, the photon flux produced from scalaron decays can constrain the decay rate from the observation of an excess photon flux over known astrophysical backgrounds. In this case, gamma-ray observations with SPI, a spectrometer telescope aboard the INTEGRAL satellite, operating in the energy range of 18 keV to 8 MeV Winkler:2003nn will be most sensitive. The photon signal from non-relativistic decays of scalaron in the present epoch yields a monochromatic spectrum, peaking at a single energy bin around, $m/2$. Consequently, the scalaron mass is constrained from the INTEGRAL/SPI data, unlike scenarios with axion-like particles (ALPs) with an undetermined U(1) symmetry breaking scale involved Alonso-Alvarez:2019ssa ; Arias:2012az . Nevertheless, there are existing limits on the decay rate of light DM decaying to two photons Calore:2022pks ; Fischer:2022pse , which can be used to constrain our scalaron scenario. In Fig.4, we show the exclusion region on the decay rate (green shaded region) using the INTEGRAL/SPI data, taken from Ref.Calore:2022pks . Authors of Ref.Calore:2022pks analyzed 16 years of the INTEGRAL/SPI data, taking the NFW profile of galactic DM density to constrain the decay rate at 95% C. L. We compare the decay rate of the scalaron (shown in red solid line) as derived in Eq.43 to find the upper bound of the scalaron mass as $m=0.7$ MeV. This bound is derived with the assumption that the scalaron satisfies total DM density, otherwise the exclusion limits become weaker. This also ensures that the scalaron can not decay to a electron-positron pair, because the mass required for that is already excluded from the excess photon flux limit.
4.3 Fifth-force constraints
The scalaron can mediate an attractive fifth force between two non-relativistic particles with masses $m_{a}$ and $m_{b}$. The resulting Yukawa-type potential takes the form Kapner:2006si ; Adelberger:2006dh
$V_{ab}(r)=-\frac{3\alpha}{16\pi M_{P}^{2}}\,\frac{m_{a}m_{b}}{r}\,e^{-mr},$ (44)
where $m$ is the scalaron mass and the dimensionless strength parameter is fixed to $\alpha=1/3$ in $R^{2}$ gravity Stelle:1977ry .
Precision torsion-balance experiments searching for deviations from Newtonian gravity have not observed any additional Yukawa-type interaction with gravitational strength ($\alpha\sim\mathcal{O}(1)$) for force ranges exceeding $38.6~\mu\mathrm{m}$ at $95\%$ C. L., as reported in Ref. Lee:2020zjt . Interpreting these null results in terms of the above Yukawa potential with $\alpha=1/3$ leads to a lower bound on the scalaron mass Lee:2020zjt ; Kapner:2006si ; Adelberger:2006dh ; Cembranos:2008gj
$m\gtrsim 2.7~\mathrm{meV}\qquad(95\%~\text{C.L.}).$ (45)
We emphasize that this constraint is specific to $R^{2}$ gravity, where the scalaron couples universally to matter with a fixed strength $\alpha=1/3$. For a generic ultralight scalar field, such as an axion-like particle, the Yukawa coupling strength $\alpha$ is a free parameter and can be arbitrarily small. In such cases, fifth-force constraints gets significantly weakened or entirely evaded by suppressing the coupling to matter Calmet:2020rpx ; Klimchitskaya:2020cnr ; Hees:2018fpg .
Figure 5: Summary of constraints on $m$ and $\xi$: The relic density constraint in scenario-I (non-zero scalaron-Higgs interaction) is shown by the red solid line, whereas that in scenario-II (negligible scalaron-Higgs interaction) is shown by the black dot-dashed line. Constraints from the LHC, the torsion balance experiment and the INTEGRAL/SPI telescope are shown using purple dashed, blue dashed and green dashed lines respectively. These constraints allow scalaron mass ranging from $2.7$ meV to $0.7$ MeV and non-minimal coupling $0\leq\xi\lesssim 10^{26}$.
5 Summary of results
[…]
Figure 6: Scalaron as dark matter in scenario-II : Constraints on the scalaron mass and its initial field value from the observed DM relic density, fifth force constraint, Higgs mass measurement and INTEGRAL/SPI limits on the excess photon flux from the scalaron decay into a pair of photons.
The situation is markedly different for scenario-II. In this case, the INTEGRAL/SPI bound becomes important and determines the upper limit on the scalaron mass, as illustrated by the green shaded region of Fig. 6, while the lower limit is set by the fifth-force constraint, shown by red shaded region in Fig. 6. Scenario-II also remains consistent with the LHC constraint, as evident from the contour defined by $3\xi m^{2}=\lambda v^{2}$, shown by the black dot-dashed line in Fig. 5. For each point on the line, the relic density is satisfied for a given $\phi_{i}$. This is depicted in Fig.6 by a constant contour (blue solid line) of $\Omega_{DM}h^{2}=0.12$. The region above the blue solid line is excluded, as the scalaron would over-saturate the DM density, while the lower region corresponds to the scalaron contributing only a fraction of the DM abundance. We note that the parameter space of the scalaron mass is substantially expanded in scenario-II with the introduction of $\xi$.
To note, the LHC bound becomes insignificant, unlike scenario-I. This can be seen explicitly by solving $3\xi m^{2}=\lambda v^{2}$ together with $|\xi m|=2.35\times 10^{16}~\mathrm{GeV}$, corresponding to the intersection of the black dot-dashed and purple dashed lines in Fig.5. This yields $m=1.1\times 10^{-13}~\mathrm{GeV}$, which is already excluded by the fifth-force constraint, apparent from the purple dashed line in Fig.6.
To summarize, the lower bound on the scalaron dark matter mass in both scenarios is robustly set by fifth-force experiments. The upper bound, however, depends primarily on the nature of the scalaron Higgs interaction. If the interaction is dominated by the Higgs non-minimal coupling to gravity, the LHC constraint determines the maximum allowed scalaron mass to be $120$ meV. Otherwise, if the Higgs quartic self-coupling dominates, the scalaron mass becomes finely tuned to approximately $3.6$ meV. Finally, in the absence of the scalaron Higgs interaction, the gamma-ray bound from the INTEGRAL/SPI telescope provides the strongest upper limit of $m\approx 0.7$ MeV.
6 Outlook
@@
In this work, we have studied the scalaron DM dynamics in presence of scalaron-Higgs trilinear interaction, which consists of Higgs quartic coupling and Higgs non-minimal coupling to gravity. Depending on the relative contributions of these couplings the scalaron dynamics gets modified and its mass is significantly altered to account for observed DM density. In particular, we provide a natural explanation for vanishing scalaron-Higgs interaction, leading to DM candidates with masses within meV-MeV.
@@
We note that the interplay between Higgs quartic coupling and the non-minimal coupling in the light DM context is demonstrated with a possibility of detection in highly sensitive X-ray/gamma ray telescopes. However, there is significant uncertainty over the experimental determination of these couplings, which can affect scalaron dynamics in the early universe. In fact, measuring the Higgs quartic coupling at colliders is very challenging, as it requires cross-section measurements of processes involving the three-Higgs vertex, which are highly suppressed Fuks:2017zkg ; Stylianou:2023tgg . On the other hand, the limits on the Higgs non-minimal coupling in the low energy effective theory with $R^{2}$ modification to the Einstein’s GR is an unexplored topic, to the best our knowledge. Here, for the first time, we have derived an upper bound on $|\xi m|$ using the Higgs mass measurement at the LHC.
Within our analysis of scenario-I, the scalaron dynamics leading to a cold DM candidate at late times is completely determined by the Higgs EWPT. We find that the Higgs non-minimal coupling can take enormously large values in contrast to Higgs-inflation paradigmBezrukov:2007ep ; Giudice:2010ka ; Rubio:2018ogq , since in the mixed model the effective scale is determined by $|\xi m|$. As long as $m$ is sufficiently small, perturbativity and the unitarity are preserved.
In contrast, within the standard misalignment framework of Scenario II, the relic density constraint is satisfied for scalaron masses ranging from $\mathcal{O}(\text{meV})$ to $\mathcal{O}(\text{MeV})$. In this regime, the scalaron dynamics can be effectively treated as decoupled from the Higgs sector for all practical purposes, with the relic abundance being primarily determined by a sufficiently large initial misalignment. As a result, the details of the electroweak crossover and the subsequent evolution of the Higgs field have a negligible impact on the scalaron evolution. This allows us to safely neglect the precise nature of the electroweak crossover in the analysis of scenario-II.
@@ Acknowledgment
The authors thank the anonymous referee for critical comments which led to significant improvement of the manuscript. SGC acknowledges funding from the Anusandhan National Research Foundation (ANRF), Govt. of India, under the National Post-Doctoral Fellowship (File no. PDF/2023/002066). SGC thanks Ananda Dasgupta, Abdul Rahaman Shaikh and Kunal Pandey for useful discussions. DG acknowledges the Post-doctoral Fellowships from IISER Kolkata and from SINP Kolkata.